# Flexible MIMO for Future Wireless Communications: Which Flexibilities are Possible?

### v1

@@ IV-B Cell-Free Movable Antenna
The cell-free network is a novel topology that can enhance the macro diversity of the network. The movable antenna technology can enhance the micro diversity of the network by fine-tuning the propagation conditions based on small-scale fading effects. Motivated by these two flexible MIMO technologies, in this case study, we combine them and study the cell-free MA technology, where antennas at each distributed AP can flexibly move along the array. Then, the sum SE maximization problem is established to optimize the antenna positions and precoding schemes. The penalty-based block coordinate descent optimization framework can be utilized to solve the above optimization problem. Fig. 5 investigates the sum SE performance for the cell-free architectures with the MA and FPA. As observed, compared with the conventional FPA-based cell-free architecture, the cell-free MA architecture can achieve much better SE performance, regardless of the centralized and distributed optimization. This case study embraces insightful motivations for the combination of all studied flexible MIMO technologies to facilitate more promising future MIMO technologies.
@@ V Conclusions
This paper presented a comprehensive exploration of the flexible MIMO technology. Twelve promising flexible MIMO technologies across three major dimensions, flexible deployment characteristics, flexible geometry characteristics, and flexible real-time modifications, were introduced. Furthermore, we identified three critical enablers: efficient CSI acquisition, low-complexity beamforming, and explainable AI-enabled optimization, along with eight sub-enabling technologies. Through two illustrative case studies on pre-optimized irregular arrays and cell-free movable antennas, we validated the substantial capacity enabled by flexible MIMO technologies.

### v2

@@ IV-B Cell-Free Movable Antenna
The cell-free network is a novel topology that can enhance the macro diversity of the system [14]. The movable antenna technology can enhance the micro diversity of the network by fine-tuning the propagation conditions based on small-scale fading effects [5]. In this case study, we combine them and creatively study the cell-free MA technology, where antennas at each distributed AP can flexibly move along the array. The sum SE maximization problem is established to jointly optimize the antenna positions and precoding schemes. The penalty-based block coordinate descent optimization framework, tailored for the cell-free movable antenna network, can be utilized to solve this optimization problem. Fig. 5 investigates the sum SE performance for the cell-free architectures with the MA and FPA. As observed, compared with the conventional FPA-based cell-free architecture, the cell-free MA architecture can achieve much better SE performance, regardless of the centralized and distributed optimization. This case study embraces insightful motivations for the combination of all studied flexible MIMO technologies to facilitate more promising future MIMO technologies. However, although antenna movement can benefit SE, the additional cost, such as motor energy and algorithm complexity, should also be considered, where a trade-off between the performance and cost should be further studied in future. Meanwhile, our case studies showcase the powerful integration of different flexible MIMO technologies, which demonstrate the synergistic and flexible evolution of future MIMO technologies we advocate in this paper.
@@ V Conclusions
This paper presented a comprehensive exploration of the flexible MIMO technology. Twelve promising flexible MIMO technologies across three major dimensions, flexible deployment characteristics, flexible geometry characteristics, and flexible real-time modifications, were introduced. Furthermore, we identified three critical enablers: efficient CSI acquisition, low-complexity beamforming, and explainable AI-enabled optimization, along with eight sub-enabling technologies. Through two illustrative case studies on pre-optimized irregular arrays-enabled high-speed railway network and cell-free movable antennas, we validated the substantial capacity enabled by flexible MIMO.